Revaluing and Contributing Assets to a Partnership

Demarco Lee invested $74,000 in the Camden & Sayler partnership for ownership equity of $74,000. Prior to the investment, equipment was revalued to a market value of $410,000 from a book value of $314,000. Kevin Camden and Chloe Sayler share net income in a 1:3 ratio.

Required:

a. Provide the journal entry for the revaluation of equipment.

For a compound transaction, if an amount box does not require an entry, leave it blank.

b. Provide the journal entry to admit Lee.
Answer #1
a
Equipment 96000 =410000-314000
      Kevin Camden, Capital 24000 =96000*1/4
      Chloe Sayler, Capital 72000 =96000*3/4
b
Cash 74000
    Demarco Lee, Capital 74000